OpenStudy (anonymous):

A ship sends out an ultrasonic sound wave with a wavelength of 5cm and a frequency of 6 x 10^4 Hz to determine the depth of the seabed. If the time interval between the transmission and detection of the reflected ultrasonic sound wave is 0.8s, what is the depth of the seabed? Explain the working please.

OpenStudy (anonymous):

Let's calculate the speed first. Speed = wavelength * frequency = 6 * 10^4 * 5 * 10^(-2) = 30 * 10^2 = 3000m/s

OpenStudy (anonymous):

Since, the distance traveled by the ultrasonic sound wave is same in both cases. Let the time taken by the wave to reach the sea-bed be "x" seconds. So, time taken by the wave to return to the ship is "0.8 - x" seconds. Now, 3000*x = 3000* (0.8 - x) Or, 2x = 0.8 Or, x= 0.4 seconds Now, depth = speed * time = 3000 m/s * 0.4s = 1200 m